what is a density of a certain material whose mass is 172g and displaxes 20ml of water ?
1
Expert's answer
2019-09-02T09:26:49-0400

We can find the density of a certain material from the formula:


"\rho = \dfrac{m}{V},"

here, "m" is the mass of the material, "V = V_{displaced}" is the volume of the material (which is equal to the volume of displaced water).

Then, we get:


"\rho = \dfrac{m}{V} = \dfrac{172g \cdot \dfrac{1 kg}{1000 g}}{20 mL \cdot \dfrac{1L}{1000mL} \cdot \dfrac{1m^3}{1000L}} = 8600 \dfrac{kg}{m^3}."

Answer:

"\rho = 8600 \dfrac{kg}{m^3}."